OpenStudy (anonymous):

what is s10 for 1+4+16+64+......?

OpenStudy (anonymous):

hai ganeshie

ganeshie8 (ganeshie8):

hey!

OpenStudy (anonymous):

help? :P can i also have steps so i know wit I'm doing?

ganeshie8 (ganeshie8):

lets see, do you notice any pattern in the given series ?

OpenStudy (anonymous):

ya multiply 4

ganeshie8 (ganeshie8):

Right, \(S_{10}\) means the sum of first \(10\) terms : \[S_{10} = 1 + 4 + 4^2 + 4^3+\cdots + 4^8 + 4^9\] just check if we have exactly \(10\) terms in above sum

OpenStudy (anonymous):

._.

OpenStudy (anonymous):

so is there a formula or something? cause this is way too long

ganeshie8 (ganeshie8):

Need more than two dots and a hyphen to continue assist you .-.

OpenStudy (anonymous):

._. thats all i got

OpenStudy (anonymous):

BRUH U DASHED OFF WHEN U COULDNT HELP ME?!!

OpenStudy (anonymous):

help :(

OpenStudy (butterflydreamer):

note: This is a geometric progression :) So, if you want to find the sum of the first 10 terms, you would use this formula: \[S_n = \frac{ a(1-r)^n }{ 1-r }\] n = nth term, r = common ratio, a = first term

OpenStudy (anonymous):

Geometric progress question...

OpenStudy (butterflydreamer):

oops i typed it wrong. Give me a minute

OpenStudy (butterflydreamer):

*\[S_n = \frac{ a(1-r^n) }{ 1-r }\]

OpenStudy (butterflydreamer):

that is the sum formula LOL

OpenStudy (anonymous):

\[a_n=a_1(r ^{n-1})\]

OpenStudy (anonymous):

so if i use this i get \[S10=\frac{ 1(1-4^2) }{ 1-4 }\]

OpenStudy (anonymous):

5 and it doesnt make sense?

OpenStudy (anonymous):

?

OpenStudy (butterflydreamer):

how did you get the 4^2 part?

OpenStudy (anonymous):

he put it inside the parentheses instead of outside

OpenStudy (anonymous):

4^10

OpenStudy (anonymous):

SRRY

OpenStudy (anonymous):

caps

OpenStudy (butterflydreamer):

yes it should be 4^10 :) Nowwww, simplify

OpenStudy (anonymous):

349525?

OpenStudy (butterflydreamer):

excellent :)!

OpenStudy (anonymous):

yea...i initial thought it was asking for the 10th term lool

OpenStudy (anonymous):

YAY! THNX :3

OpenStudy (anonymous):

until i saw its S not a

OpenStudy (butterflydreamer):

no worriesss ^_^!